# Quasi-invariant states with uniformly bounded cocycles

### v2

@@ 4 Unitary transformations associated to quasi–invariant states
In the setting of Theorem 3.2, we have that there exists an invertible $a\in\mathcal{A}$ such that $D_{\psi}=D_{\varphi}^{1/2}a^{2}D_{\varphi}^{1/2}$. Thus there is a partial unitary $v$ such that $D_{\psi}^{a/2}=vaD_{\varphi}^{1/2}$. Since both operators have full support, $v$ is a unitary and we deduce that there exists an invertible $\gamma\in\mathcal{A}$ such that $D_{\psi}^{1/2}=\gamma D_{\varphi}^{1/2}$. Thus we get $D_{\psi}=dD_{\varphi}=\gamma D_{\varphi}\gamma^{*}$. As above, we get that $\gamma^{*}$ sits in the domain of $\sigma^{\varphi}_{-i}$ and $\sigma^{\varphi}_{-i}(\gamma^{*})=\gamma d$. We conclude that $d=\gamma^{-1}\sigma^{\varphi}_{-i}(\gamma^{*})$.
In the same way, let $\gamma_{g}=\gamma^{-1}g^{-1}(\gamma)$ for $g\in G$. then one can also check that $\gamma_{g}^{*}$ sits in the domain of $\sigma^{\varphi}_{-i}$ and that $x_{g}=\gamma_{g}^{-1}\sigma^{\varphi}_{-i}(\gamma_{g}^{*})$.
@@
If $\varphi$ is $G$-strongly quasi-invariant then $g\mapsto U_{g}$ is a unitary representation of the group $G$.
Proof. Notice that if $\varphi$ is $G$-strongly quasi-invariant, it is showed in [2] that the cocycles $x_{g},\;g\in G$, are in the centralizer of $\varphi$. Therefore by (4.3) it follows that $a^{2}_{g}=x_{g}=x_{g}^{*}$ (i.e. $a_{g}=x_{g}^{1/2}$). Moreover from (4.2), one has
@@
By using the cocycle property it is easy to check that $U_{g}U_{h}=U_{gh}$ for all $g,h\in G$ and thus $g\mapsto U_{g}$ is a representation of $G$, which satisfies $U_{g}^{*}=U_{g}^{-1}$. $\Box$
5 A Theorem of Kovács and Szücs for quasi–invariant states with uniformly bounded cocycles
@@
Since $\psi$ is a normal faithful $G$-invariant state on $\mathcal{A}$, it follows from [9, Theorem 1] that there exists a normal faithful $G$-invariant Umegaki conditional expectation $\Phi:\mathcal{A}\rightarrow\mathcal{B}$, where $\mathcal{B}$ is a sub-von Neumann algebra of $\mathcal{A}$, such that
@@
Now, if $\varphi$ is a $G$-strongly quasi-invariant state and if $g\mapsto U_{g}$ is the unitary representation of $G$ defined by Theorem 4.2, we will show that the conditional expectation $\Phi$ in (5.1) can also be characterized by $U_{g},\;g\in G$ Define
@@
Let $E_{0}$ be the orthogonal projection onto the subspace of all $\xi\in L_{2}(\mathcal{A})$ invariant under all $U_{g}$. We denote by $\mathcal{B}$ the set of fixed points of $\mathcal{A}$ under the action of the automorphisms $u_{g}$, $g\in G$. Let $F_{0}=[\mathcal{B}^{\prime}E_{0}]$ be the orthogonal projection onto $\overline{\mbox{span}(\mathcal{B}^{\prime}E_{0}L_{2}(\mathcal{A}))}$. Finally, let $\mathcal{M}$ be a Godement mean over $G$. The following result follows from a theorem of Kovács and Szücs (see [14] or [1] and [9, Theorem 1]).
@@
    There exists a unique normal faithful $G$-invariant Umegaki conditional expectation $\Phi:\mathcal{A}\rightarrow\mathcal{B}$.
    Moreover, $\Phi(b)$ can be defined (equivalently) as
@@
      the unique element of $\mathcal{A}$ such that Φ(b)E0

### v3

@@ 4 Unitary implementation associated to quasi–invariant states
In the setting of Theorem 3.2, we have that there exists an invertible $a\in\mathcal{A}$ such that $D_{\psi}=D_{\varphi}^{1/2}a^{2}D_{\varphi}^{1/2}$. Thus there is a partial isometry $v$ such that $D_{\psi}^{1/2}=vaD_{\varphi}^{1/2}$. Since both operators have full support, $v$ is unitary and we deduce that there exists an invertible $\gamma\in\mathcal{A}$ such that $D_{\psi}^{1/2}=\gamma D_{\varphi}^{1/2}$. Thus we get $D_{\psi}=d^{*}D_{\varphi}=\gamma D_{\varphi}\gamma^{*}$. As above, we get that $\gamma^{*}$ sits in the domain of $\sigma^{\varphi}_{-i}$ and $\sigma^{\varphi}_{-i}(\gamma^{*})=\gamma^{-1}d^{*}$. We conclude that $d^{*}=\gamma\sigma^{\varphi}_{-i}(\gamma^{*})$. In the same way, let $\gamma_{g}=g^{-1}(\gamma^{-1})\gamma$ for $g\in G$, then one can also check that $\gamma_{g}^{*}$ sits in the domain of $\sigma^{\varphi}_{-i}$ and that $x_{g}^{*}=\gamma_{g}\sigma^{\varphi}_{-i}(\gamma_{g}^{*})$.
@@
If $\varphi$ is $G$-strongly quasi-invariant then $g\mapsto U_{g}^{*}$ is a unitary representation of the group $G$.
Proof. Notice that if $\varphi$ is $G$-strongly quasi-invariant, it is showed in [2] that the cocycle $x_{g},\;g\in G$, are in the centralizer of $\varphi$. Therefore by (4.3) it follows that $a^{2}_{g}=x_{g}=x_{g}^{*}$ (i.e. $a_{g}=x_{g}^{1/2}$). Moreover from (4.2), one has
@@
By using the cocycle property, for any $g,h\in G$, $x\in\mathcal{A}$, one has
$U_{g}U_{h}(xD_{\varphi}^{1/2})=U_{g}(h^{-1}(x)x_{h}^{1/2}D_{\varphi}^{1/2})=g^{-1}(h^{-1}(x))g^{-1}(x_{h}^{1/2})x_{g}^{1/2}D_{\varphi}^{1/2}=U_{hg}(xD_{\varphi}^{1/2}).$
It follows that $g\mapsto U_{g}^{*}=U_{g^{-1}}$ is a unitary representation of the group $G$. $\Box$
5 A Theorem of Kovács and Szücs for quasi–invariant states with uniformly bounded cocycle
@@
Since $\psi$ is a normal faithful $G$-invariant state on $\mathcal{A}$, it follows from [8, Theorem 1] that there exists a normal faithful $G$-invariant Umegaki conditional expectation $\Phi:\mathcal{A}\rightarrow\mathcal{B}$, where $\mathcal{B}$ is a sub-von Neumann algebra of $\mathcal{A}$, such that
@@
Now, if $\varphi$ is a $G$-strongly quasi-invariant state and if $g\mapsto U_{g}^{*}$ is the unitary representation of $G$ defined by Theorem 4.2, we will show that the conditional expectation $\Phi$ in (5.1) can also be characterized by $U_{g},\;g\in G$ Define
@@
Let $E_{0}$ be the orthogonal projection onto the subspace of all $\xi\in L_{2}(\mathcal{A})$ invariant under all $U_{g}$. We denote by $\mathcal{B}$ the set of fixed points of $\mathcal{A}$ under the action of the automorphisms $u_{g}$, $g\in G$. Let $F_{0}=[\mathcal{B}^{\prime}E_{0}]$ be the orthogonal projection onto $\overline{\mbox{span}(\mathcal{B}^{\prime}E_{0}L_{2}(\mathcal{A}))}$. Finally, let $\mathcal{M}$ be a Godement mean over $G$. The following result follows from a theorem of Kovács and Szücs (see [14] or [1] and [8, Theorem 1]).
@@
    There exists a unique normal faithful $G$-invariant conditional expectation
    $\Phi:\mathcal{A}\rightarrow\mathcal{B}$. Moreover, $\Phi(b)$ can be defined (equivalently) as
@@
      the unique element of $\mathcal{A}$ such that $\Phi(b)E_{0}=E_{0}bE_{0}$
    2. (ii)
      the unique element of $\mathcal{A}$ such that
      $(\xi_{1},\;\Phi(b)\xi_{2})=\mathcal{M}\{(\xi_{1},\;U_{g^{-1}}bU_{g}\xi_{2})\}$
  2. (2)
    A faithful normal state $w$ on $\mathcal{A}$ is $G$-strongly quasi-invariant with uniformly bounded cocycle if and only if there exists a positive invertible bounded operator $d$ such that $w(.)=(w(d.)|\mathcal{B})\circ\Phi(d^{-1}.)$
Proof. Let $\varphi$ is a faithful normal $G$-strongly quasi-invariant state with uniformly bounded cocycle. By applying Theorem 3.3, there exists a bounded positive invertible operator $d$ and a state a $G$-invariant state $\psi$ such that
$\psi(a)=\varphi(da)=\langle d^{1/2}D_{\varphi}^{1/2},ad^{1/2}D_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})}\quad\forall a\in\mathcal{A}.$
Remember that $x_{g}=dg^{-1}(d^{-1})=x_{g}^{*}$. Moreover for all $g\in G$
$U_{g}d^{1/2}D_{\varphi}^{1/2}=g^{-1}(d^{1/2})x_{g}^{1/2}D_{\varphi}^{1/2}=d^{1/2}D_{\varphi}^{1/2}.$
Let $p$ be the orthogonal projection onto $\overline{\mbox{span}(\mathcal{B}^{\prime}d^{1/2}D_{\varphi}^{1/2}})$. Then $p\in\mathcal{B}$ and $p\leq F_{0}$. Moreover, one has
$\psi(p)$ $=$ $\langle d^{1/2}D_{\varphi}^{1/2},pd^{1/2}D_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})}$
$=$ $\langle d^{1/2}D_{\varphi}^{1/2},d^{1/2}D_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})}$
$=$ $\varphi(d)=\psi(I)=1.$
Since $\psi$ is faithful, $p=I$ and $F_{0}=I$. Finally the results of the above theorem follows from [8, Theorem 1]. $\Box$
Now we provide a commutative easy example of faithful normal $G$-quasi-invariant state such that the cocycle is not uniformly bounded. Moreover, we show in this example that there is no $G$-invariant conditional expectation.
Example. On $\mathcal{A}=L^{\infty}(\mathbb{R})$, define the state
[…]
The additive group $\mathbb{R}$ acts on $\mathcal{A}$ by translation:
$\tau_{t}(f)(s):=f(s-t)$
It is clear that $\varphi$ is $\mathbb{R}$-quasi-invariant state with bounded cocycle
$x_{t}(s)=\frac{1+s^{2}}{1+(s+t)^{2}}.$
The cocycle $(x_{t})_{t}$ does not satisfy assumption (A) since $\|x_{t}\|,\|x_{t}^{-1}\|\geq 1+t^{2}$. The fixed point of the action are constant functions. Thus every conditional expectation $\Phi$ on $L^{\infty}(\mathbb{R})$ has to be of the form
$\Phi(f)=\psi(f)1$ (5.3)
for some normal state $\psi$. But there are no functions in $L_{1}$ that are $\mathbb{R}$-invariant.
This example also shows that it seems unlikely to improve Theorem 3.2.
6 Strongly quasi-invariant states and semifinite trace
In this section we extend the result developped by $\o$rmer [20] to the case of strongly quasi-invariant states with uniformly bounded cocycles. Given a semifinite von Neumann algebra $\mathcal{A}$, a group $G$ of *-automorphisms is said to act ergodically on the center $\mathcal{Z}=\mathcal{A}\cap\mathcal{A}^{\prime}$ of $\mathcal{A}$ if $\mathcal{F}(G)\cap\mathcal{Z}=\mathbb{C}1$, where $\mathcal{F}(G)$ is the set of the fixed points of $G$ in $\mathcal{A}$.
Theorem 6.1
Let $\mathcal{A}$ be a semifinite von Neumann algebra and let $G$ be an amenable group of *-automorphisms of $\mathcal{A}$ and assume that $G$ acts ergodically on the center $\mathcal{Z}$ of $\mathcal{A}$. Suppose that $\varphi$ is a faithful normal $G$-quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}$’s satisfying assumption $(A)$. Then, there exists up to a scalar multiple a unique faithful normal $G$-invariant semifinite trace $\tau$ of $\mathcal{A}$. Furthermore the density $c\in L_{1}(\mathcal{A},\tau)^{+}$ of $\varphi$ with respect to $\tau$ satisfies
$(g^{-1})^{*}(c)=cx_{g^{-1}},\quad x_{g}^{*}c=cx_{g},\quad\forall g\in G,$ (6.1)
where $(g^{-1})^{*}:L_{1}(\mathcal{A},\tau)\rightarrow L_{1}(\mathcal{A},\tau)$ is the predual map of $g^{-1}$ on $\mathcal{A}$.
Proof. Under assumptions of the above theorem, it follows from Theorem 3.2 that there exists a normal faithful $G$-invariant state $\psi$. Applying [[20], Theorem 1] for $\psi$, there exists up to a scalar multiple a unique faithful normal $G$-invariant semifinite trace $\tau$ on $\mathcal{A}$. Let $c\in L_{1}(\mathcal{A},\tau)^{+}$ be the density of $\varphi$. For all $a\in\mathcal{A}$, we have
$\tau((g^{-1})^{*}(c)a)=\tau(cg^{-1}(a))=\varphi(g^{-1}(a))=\varphi(x_{g^{-1}}a)=\tau(cx_{g^{-1}}a).$
This proves the first equality (6.1). For the second, note that for all $a\in\mathcal{A}$
$\tau(cx_{g}a)=\varphi\circ g(a)=\overline{\varphi\circ g(a^{*})}=\overline{\tau(cx_{g}a^{*})}=\tau(x_{g}^{*}ca).$
This implies that $x_{g}^{*}c=cx_{g}$ in $L_{1}(\mathcal{A},\tau)$. $\Box$
Example. Taking again $\mathcal{A}=L^{\infty}(\mathbb{R})$ and $G$ to be the $ax+b$ group of all automorphisms of $\mathcal{A}$ of the form $\pi_{(a,b)}f(t)=f(at+b)$ for $(a,b)\in\mathbb{R}^{*}\times\mathbb{R}$. Then $G$ is amenable (since it is solvable), it admits a cocycle for the state given by the measure $\frac{1}{\pi}\frac{1}{1+t^{2}}dt$; $x_{(a,b)}(t)=\frac{a(1+t^{2})}{a^{2}+(t-b)^{2}}$ that is not uniformly bounded. Moreover, one can check that the fixed point algebra by $G$ consists only in constant functions. Similarly, one also easily sees that there are no measures on $\mathbb{R}$ invariant by $G$. Thus one cannot remove the assumption $(A)$ even in a commutative situation.
Acknowledgement
A. Dhahri is a member of GNAMPA-INdAM and he has been supported by the MUR grant Dipartimento di Eccellenza 2023-2027 of Dipartimento di Matematica, Politecnico di Milano.
Statements and declarations
Ethical Approval: Not applicable.
Competing interests: The authors declare that they have no conflict of interest.
Data availability statement: The manuscript has no associated data.